Jane and some students from her geometry class went outside to measure their shadows. Jane measured Tonya's and Seth's shadows and recorded the information in the table below
Height(x) shadow length(y)
158cm 106cm
166cm 112cm
Using the information in the table provided above and answer the following and show your work:
a. Find the slope of a line representing this data in the table above.
b. Using the slope and one of the points in the table, write an equation in slope-intercept form representing the student's shadow in relation to height.
c. Jane forgot to include her shadow in the student data. She measured her height and recorded 164 cm. What should she record for her shadow length?

Answer by stanbon(75887)

Jane and some students from her geometry class went outside to measure their shadows. Jane measured Tonya's and Seth's shadows and recorded the information in the table below
Height(x) shadow length(y)
158cm 106cm
166cm 112cm
Using the information in the table provided above and answer the following and show your work:
a. Find the slope of a line representing this data in the table above.
slope = (112-106)/(165-158) = 6/8= 3/4
--------------------------------------
b. Using the slope and one of the points in the table, write an equation in slope-intercept form representing the student's shadow in relation to height.
shadow = (3/4)height + b
Substitute to solve for "b":
106 = (3/4)158+b
b = 106-118.5
b = -12.5
EQUATION:
s(h)=(3/4)h-12.5
----------------------
c. Jane forgot to include her shadow in the student data. She measured her height and recorded 164 cm. What should she record for her shadow length?
s(164) = (3/4)164-12.5
s(164) = 110.5 cm
--------------
Cheers,
Stan H.